# UDC-VIT: A Real-World Video Dataset for Under-Display Cameras

### v1

@@ A.1 Dataset Acquisition
The alignment algorithm we use involves shifting, rotating, and cropping paired frames with DFT. The detailed alignment algorithm is illustrated in Algorithm A.1. In this algorithm, following Ahn et al. (2024)’s setting, we use $\lambda_{1}=\lambda_{3}=1$ and $\lambda_{2}=0$, and we do not apply rotation. Their experiments show that applying rotation reduces the Percentage of Correct Keypoints (PCK) when varying $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\theta_{\text{rotation}}$.
The loss function in Equation 1 in the main body enables the incorporation of both local (i.e., MSE) and global (i.e., DFT) information across spatial and frequency domains. Using DFT to align the paired frames offers a significant advantage because it can decompose a frame into its constituent spatial frequency components. Figure A.1(a) and (c) depict paired frames $\mathcal{G}$ and $\mathcal{D}$ comprising multiple sinusoidal gratings, indicating a noticeable spatial shift. Figure A.1(b) and (d) represent the phase and amplitude differences between the paired frames, respectively. Thus, reducing the phase component is critical for effectively aligning the paired frames for the same scene.
Algorithm A.1 Alignment of paired images $I_{G}$ and $I_{D}$ (Ahn et al., 2024).
Images $I_{G}$, $I_{D}$ of size $(H,W)$, hyperparameters $s$, $\theta_{r}$, $r$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$
@@
     for $x_{\text{shift}}$ from $-s$ to $s$ with step 1 do
         for $y_{\text{shift}}$ from $-s$ to $s$ with step 1 do
@@
              Crop image $\mathcal{D}_{\text{tmp}}$ from $\mathcal{D}_{\text{rotated}}$ at position $(p,q)$
@@
Figure A.1: Frequency analysis based on the conceptual illustration for paired frames involving shifts without degradation. (a) The original frame $\mathcal{G}$ consists of multiple sinusoidal gratings. The inverse $DFT$ applied to $\mathcal{F}_{G}(u,v)$ produces each sinusoidal grating. (b) The phase difference between $\mathcal{G}$ and $\mathcal{D}$. (c) The spatially shifted frame $\mathcal{D}$ in the spatial domain comprises multiple sinusoidal gratings, as in (a). (d) The amplitude difference between $\mathcal{G}$ and $\mathcal{D}$, showing no difference.
The controller.
When capturing videos, we discard the initial 30 frames because it takes approximately 15 frames for the ground-truth camera and 25 frames for the UDC to achieve focus. The UDC requires more frames for focusing due to its degradation. Furthermore, we use a solid-state drive (SSD) instead of a secured digital (SD) card, since the SD card takes longer to save FHD resolution videos, which disrupts synchronization between the two cameras.
A.2 Dataset Details and Statistics
@@
From a pool of 647 videos, we have randomly selected 510 for training, 69 for validation, and 68 for the test set. The UDC-VIT dataset is available in PNG format accompanied by a conversion script from PNG to NPY. We have annotated each video pair, providing a detailed overview of the total count and the distribution of different annotation labels. The video pairs are thoughtfully categorized into various settings, including the presence of flare and light sources, human presence and types of human motion, and indoor/outdoor.
Figure A.2: The dataset distribution. The parenthesis beside a label is the encoding of the label. Note that a video pair can have multiple annotation labels. The distribution of (a) lighting conditions, (b) human presence and their actions, and (c) shooting location.
Table A.1: Prescribed instructions from IRB-approved participant information sheet. Out of thirty shots per person, videos displaying issues such as being out-of-focus are eliminated from the dataset.
@@
A. If the participants agree to take part in, the following procedures will be conducted:
The participants will be photographed with 30 shots using the UDC and regular digital cameras according to the following motions:
@@ A. If the participants agree to take part in, the following procedures will be conducted:
Since the UDC is located under the display and operates in low-light environments, it is necessary to shoot in various locations (indoors/outdoors) and lighting conditions (bright/dark) account for the varied degradation patterns of the UDC. Additionally, recognizing individuals from different angles is essential for tasks like face recognition, particularly in financial authentication. This necessitates deep-learning models capable of restoring a subject’s appearance from different perspectives (e.g., front, left, and right), requiring a dataset with multi-angle captures. The recorded videos will be publicly released as a dataset for the UDC research.
@@
A. As a token of appreciation for their participation, participants will receive a Starbucks gift card worth 50,000 Korean won. However, if they withdraw before completing all 30 shots or request the deletion of the captured videos, we regret to inform them that compensation cannot be provided. Compensation will only be given to those who successfully complete the full 30-shot video capture. If participants request video deletion after receiving compensation, they will be required to return the full amount.
@@
We have obtained the approval of the Institutional Review Board (IRB) for the UDC-VIT dataset, as our research involves human subjects. This rigorous process ensures the highest standards of research ethics. Using IRB-approved procedures, we enlisted 22 voluntary research participants. As shown in Table A.1, the IRB-approved information sheet provides comprehensive instructions to the participant.
Similarly, it is essential to note that the users of the UDC-VIT dataset are engaged in research involving human subjects. Therefore, the users are required to secure IRB approval by the regulations of their respective countries. When users download the dataset, there will be instructions on IRB approval, as shown in Figure A.3.
A.3 Rigorous Maintenance Plan
This section provides easy accessibility and a rigorous maintenance plan for the UDC-VIT dataset for long-term preservation.
@@ A.3 Rigorous Maintenance Plan
The UDC-VIT dataset will be publicly available on the website of our research group as depicted in Figure A.3. Users can access the dataset by filling out a form on the research group’s website. Upon submission, they will receive an email with the download link. Instructions for accessing the UDC-VIT dataset are available on our project site. Distributing the dataset through the research group’s website ensures long-term preservation. Handling contact and bug reports via email allows for continuous maintenance and updates.
@@
The UDC-VIT dataset is licensed under Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International (CC BY-NC-SA 4.0). Under this license, users of the UDC-VIT dataset can freely utilize, share, and modify this work by adequately attributing the original author, distributing any derived works under the same license, and utilizing it exclusively for non-commercial purposes. It is essential to mention that the UDC-VIT dataset is restricted to UDC research purposes only, as outlined in our IRB documentation. Detailed information on this license can be found in the official Creative Commons website.
Figure A.3: The UDC-VIT dataset page on our research group’s website, which provides information and a download link.
Appendix B Analyzing the Limitations of Synthetic Datasets
Figure B.1: The visual illustration that showcases improbable flares resulting from excessive scaling in the VidUDC33K dataset (Liu et al., 2024). (a) Flare in the river. (b) Flare from the dust on the camera lens. (c) Flare on the bird feathers. (d) Flare on the flower petals. (e) Flare on the mountain peaks. (f) Flare on the food. (g) Flare in the snake eyes. (h) Flare on the waterfalls.
This section provides a comprehensive analysis of the limitations of synthetic datasets (e.g., VidUDC33K (Liu et al., 2024)). As discussed in Section 4 in the main body, the VidUDC33K dataset showcases less meaningful and strange scenes. In this section, we delve into three main strange phenomena that occur within it: the flare appearance in improbable situations and unintended white artifacts, as well as the darkened and nearly featureless degraded frames.
Improbable situations and unintended white artifacts.
Liu et al. (2024) endeavor to synthesize flares through the convolution of the PSF with ground-truth images. However, the desired flares do not manifest as expected. Subsequently, they employ a scaling procedure to pixel values exceeding a certain threshold to amplify those values, which is followed by PSF convolution.
This results in two phenomena. One is the flare appearance in improbable situations. Their method creates flares with values of white pixels above a certain threshold regardless of whether it is the light source, as shown in Figure 8 in the main body and Figure B.1. The other is unintended white artifacts. Since pixel values exceeding a certain threshold are amplified, some regions become white, losing their original colors. For example, areas with clouds in the sky, waterfalls, and white walls become excessively white, as depicted in Figure B.2. We conduct an experiment to verify that the scaling is related to the flare generation. The results are presented in the final row of Figure B.2. Flares do not manifest even in frames where they are expected to appear without the scaling procedure, as shown in Figure B.2(d). Approximately 12% of the videos exhibit unintended white artifacts due to the scaling procedure, which is unsuitable for deep-learning training.
Figure B.2: The visual depiction that shows white artifacts resulting from excessive scaling in the VidUDC33K dataset (Liu et al., 2024). The frames without the scaling procedure do not exhibit these white artifacts (see (c) and (d)), unlike the frames with the scaling procedure (see (a) and (b)). The flares are visible with the scaling procedure (see (b)), while flares are not visible without the scaling procedure (see (d)). It appears that the authors use scaling to generate flares, inadvertently creating unrealistic white artifacts in the process. (a) The ground-truth frame with scaling procedure. (b) The degraded frame with scaling procedure. (c) The ground-truth frame without scaling procedure. (d) The degraded frame without scaling procedure.
The darkened and nearly featureless frames.
Liu et al. (2024) strive to create temporally variant flares in continuous video sequences. They simulate the dynamic changes of the PSF during motion by computing the inter-frame homography matrix $H_{t-1\rightarrow t}$, formulated as Equation 3, between consecutive frames.
$k_{t}$ $=\mathcal{T}(k_{t-1},H_{t-1\rightarrow t})$ (3)
$H_{t-1\rightarrow t}$ $=\mathcal{M}(I_{t-1}^{GT},I_{t}^{GT}),$
However, this process occasionally results in PSF values approaching zero, causing the degraded frames to appear entirely black. Specifically, this issue occurs in 4 out of 677 videos. The first frame does not undergo PSF transformation, while subsequent frames do. Therefore, as seen in Figure B.3(c), the frames after the first one (e.g., the tenth frame) sometimes become black.
Figure B.3: The visual representation that demonstrates black frames resulting from incorrectly transformed PSFs in the VidUDC33K dataset Liu et al. (2024). (a) The first frame of the degraded video. (b) The first frame of the ground-truth video. (c) The tenth frame of the degraded video. (d) The tenth frame of the ground-truth video.
Appendix C Comprehensive Analysis of UDC-VIT
This section provides a thorough quantitative and qualitative analysis.
C.1 Quantitative Results
Among the categories illustrated in Figure A.2, the light conditions and shooting location are related to the restoration performance. In Table C.1, although the presence of humans seems to influence restoration performance, it is not directly correlated. To ensure the safety of participants, 86.4% of scenes, including humans, are captured indoors, which causes less severe degradation than outdoor natural flares. Given that a UDC-VIT video can have multiple annotations (e.g., an outdoor scene with flares caused by natural sunlight), the annotation type listed in a column in Table C.1 cannot be considered the only factor influencing UDC degradation. However, it is reasonable to recognize the annotation type as a significant factor affecting PSNR, SSIM, and LPIPS values.
Light sources.
As shown in Table C.1, all models encounter difficulties in restoring scenes with flare (Flare - Present - Average) compared to those without flare (Flare - Absent). Within flare-present scenes, the severity of degradation varies based on the light source (e.g., natural sunlight, artificial light, or both). Intense sunlight can oversaturate pixel values, obscuring objects around the flares (see Figure C.1(a) and (c)). Consequently, the benchmark models have more difficulty restoring videos with flares caused by natural sunlight than those caused by artificial light.
Present Absent Indoor Outdoor Present Absent
Natural Artificial Both Average
sunlight light
DISCNet (Feng et al., 2021) SSIM $\uparrow$ 0.7704 0.8451 0.8191 0.8303 0.8611 0.8573 0.7852 0.8795 0.7636 0.8403
UDC-UNet (Liu et al., 2022b) SSIM $\uparrow$ 0.8197 0.8995 0.8857 0.8856 0.9027 0.9092 0.8324 0.9276 0.8198 0.8911
FastDVDNet (Tassano et al., 2020) SSIM $\uparrow$ 0.7902 0.8523 0.8244 0.8392 0.8538 0.8593 0.7940 0.8720 0.7891 0.8439
EDVR (Wang et al., 2019) SSIM $\uparrow$ 0.7711 0.8422 0.8132 0.8276 0.8445 0.8484 0.7833 0.8612 0.7781 0.8331
ESTRNN (Zhong et al., 2020) SSIM $\uparrow$ 0.7997 0.8818 0.8577 0.8662 0.8637 0.8847 0.8025 0.8938 0.8098 0.8654
DDRNet (Liu et al., 2024) SSIM $\uparrow$ 0.8015 0.8628 0.8455 0.8512 0.8594 0.8697 0.8025 0.8810 0.8007 0.8539
Shooting location.
The benchmark models struggle more with restoring outdoor scenes than indoor scenes, as shown in Table C.1. Approximately 33.3% of outdoor and 15.4% of indoor scenes include flares caused by natural sunlight in the UDC-VIT dataset. Moreover, sunlight-induced flares occurring indoors are often less severe than those occurring outdoors. For example, outdoor scenes with natural sunlight flares show severe flare, as shown in Figure C.1(a), whereas indoor scenes with the same type of flares tend to be less severe, as depicted in Figure C.1(c) and (d). Notably, the flare in the upper right corner of Figure C.1(d) is mild, a result of sunlight scattered by a glass window rather than entering the camera directly. This understanding is crucial as it highlights the unique challenges of restoring outdoor scenes where direct sunlight is a significant factor. Consequently, all models face more significant difficulties restoring outdoor scenes than indoor scenes.
C.2 Qualitative Results
This section presents the visual results of the restored frames. The restoration results from the benchmark models still exhibit various forms of degradation, demonstrating their limitations in effectively handling degradations of real-world datasets. This underscores the novelty of the UDC-VIT dataset and the need to develop deep-learning models trained on real-world datasets.
Light sources.
As illustrated in Figure C.1, flares can be categorized into glare, shimmer, and streak (Ahn et al., 2024; Dai et al., 2022). A glare is characterized by intense and robust light, resulting in circular patterns as artifacts. Shimmer entails rapid and nuanced light or color intensity variations across an image. A streak manifests as a lengthy, slender, and usually irregular line of light or color within an image.
As outlined in Section 4 in the main body, flares exhibit variations depending on the light sources (e.g., light source variant flare) and location (e.g., spatially variant flare). In Figure C.1(a) and (c), sunlight-induced flares are intense, causing all models to struggle to restore obscured objects. In contrast, flares from artificial light, shown in Figure C.1(b) and (c), are relatively easier to restore than sunlight-induced flares. However, benchmark models still face challenges restoring areas affected by shimmer and streak in flares from artificial light, resulting in speckled artifacts around the flare. The mild flare caused by natural light, depicted in Figure C.1(d), originates from sunlight scattered by a glass window, which all models restore well. Since light sources covered by a diffuser produce less severe flares, all models effectively restore them, as shown in Figure C.1(e). Interestingly, some models fail to restore the reflected light on the human face, even though they all restore the flare caused by fluorescent light, as illustrated in Figure C.1(f). Reflected light might decrease the accuracy of face recognition.
Shooting location.
The restoration performance is sometimes influenced by the presence or absence of flares within the frame rather than solely by the shooting location. For instance, while Figure C.1(b) and (h) portray indoor scenes, models generally excel in restoring the flare-free frame in Figure C.1(h). Likewise, in outdoor scenes depicted in Figure C.1(a) and (i), models tend to achieve better restoration for the flare-free frame in Figure C.1(i). However, it is worth noting that some models may inaccurately render the sky with a reddish hue.
Human.
The presence of humans alone does not pose a significant challenge to restoration. Instead, the restoration difficulty hinges on how UDC degradations, such as noise, blur, transmittance decrease, and flare, impact humans. For example, in Figure C.1(d) and (e), despite the presence of flares in the frames, they do not affect humans. However, in Figure C.1(f), the reflection of fluorescent light on the person’s glasses poses challenges to restoring fine details around the eyes. In Figure C.1(g) and (h), human faces appear reddish in the input frames compared to the ground-truth frames due to UDC-induced diffraction occurring differently across RGB channels. In addition, the restored facial colors vary among the models. Considering the importance of facial color in images and videos, accurately addressing these diverse UDC degradations is essential for human restoration in applications such as face recognition and video conferencing.
Flicker.
The visual comparison presented in this paper is limited to a single frame. While specific frames in Figure C.1 exhibit successful restoration, temporal inconsistencies, such as flickering, are frequently observed across multiple frames in the video for all models. This flickering may result from varying degradations between consecutive frames, such as decreases in transmittance and spatially and temporally variant flares. To see the flickering of the restored videos, please visit our project site.
Appendix D Cross-dataset Validation
This section demonstrates the cross-dataset validation to tackle the unique dataset distribution and degradation patterns of UDC datasets as discussed in Section 6. For example, Samsung Galaxy Z-Fold 3 (Samsung Electronics Co., Ltd., 2021) (UDC-SIT (Ahn et al., 2024)) and Samsung Galaxy Z-Fold 5 (Samsung Electronics Co., Ltd., 2023) (UDC-VIT) share similar pixel designs, they still exhibit differences. Similarly, Samsung Galaxy Z-Fold 5 (UDC-VIT) and ZTE Axon 20 (ZTE Corporation, 2020) (VidUDC33K (Liu et al., 2024)) have vastly different pixel designs, as they come from different vendors.
Figure D.1(a) and (c) illustrate that the UDC-SIT and UDC-VIT datasets show similar degradation, such as blur, decrease in transmittance, and flare shape. In contrast, Figure D.1(b) and (c) highlight the stark difference between the VidUDC33K and UDC-VIT datasets. This discrepancy arises from two factors: the variation in pixel design and the synthetic nature of the VidUDC33K dataset, which results in unrealistic degradation patterns.
Fine-tuning models to address variant dataset distributions or degradation patterns is crucial in practical applications. To evaluate the effect of fine-tuning and validate the effectiveness of UDC-VIT, which reflects real-world degradation, we conduct three experiments (Exp. 1-3), as shown in Table D.1. The subscripts beside the model name $\mathcal{M}$ specify the datasets used for training and fine-tuning. For example, $\mathcal{M}_{s3}$ refers to the model trained on UDC-SIT (Samsung Galaxy Z-Fold 3) without fine-tuning, $\mathcal{M}_{s5z20}$ is trained on UDC-VIT (Samsung Galaxy Z-Fold 5) and fine-tuned on VidUDC33K (ZTE Axon 20), while $\mathcal{M}_{z20s5}$ is trained on VidUDC33K (ZTE Axon 20) and fine-tuned on UDC-VIT (or Samsung Galaxy Z-Fold 5). We use models $\mathcal{M}$ such as UDC-UNet (Liu et al., 2022b), DISCNet (Feng et al., 2021), and DDRNet (Liu et al., 2024) among six benchmark models in Table 3. Fine-tuning is performed for 10% or 20% of the total iterations, with the learning rate set to 10% or 20% of the original value.
Figure D.1: Comparison of the UDC datasets, showing varied data distribution and degradation patterns. (a) UDC-SIT (Ahn et al., 2024). (b) VidUDC33K (Liu et al., 2024). (c) UDC-VIT.
Table D.1: The design of experiments to verify the effect of fine-tuning and the the use of a real-world dataset (e.g., UDC-VIT). The first and the second subscripts beside $\mathcal{M}$ indicate the training and fine-tuning datasets, respectively. For example, $\mathcal{M}_{s3}$ refers to the model trained on UDC-SIT without fine-tuning, while $\mathcal{M}_{s3s5}$ denotes the model trained on UDC-SIT and subsequently fine-tuned on UDC-VIT. Models without subscripts are trained and tested on the same dataset.
Exp. 1 $\mathcal{M}_{s3}$ UDC-SIT - UDC-VIT
$\mathcal{M}_{s3s5}$ UDC-SIT UDC-VIT UDC-VIT
$\mathcal{M}$ UDC-VIT - UDC-VIT
Exp. 2 $\mathcal{M}_{s5}$ UDC-VIT - VidUDC33K
$\mathcal{M}_{s5z20}$ UDC-VIT VidUDC33K VidUDC33K
$\mathcal{M}$ VidUDC33K - VidUDC33K
Exp. 3 $\mathcal{M}_{z20}$ VidUDC33K - UDC-VIT
$\mathcal{M}_{z20s5}$ VidUDC33K UDC-VIT UDC-VIT
$\mathcal{M}$ UDC-VIT - UDC-VIT
Experiment 1: impact of fine-tuning on UDC-VIT.
[…]
These findings lead to the following conclusions. Although Samsung Galaxy Z-Fold 3 (UDC-SIT) and Samsung Galaxy Z-Fold 5 (UDC-VIT) share similar pixel designs because they are from the same vendor, their differences are substantial enough to require fine-tuning. With adequate adaptation, however, these models effectively leverage degradations from other UDC devices, underscoring the potential for cross-device generalization with fine-tuning.
Table D.2: [Exp. 1] Restoration performance of DISCNet (Feng et al., 2021) and UDC-UNet (Liu et al., 2022b) when tested on UDC-VIT. They are trained on UDC-SIT (Ahn et al., 2024), either with or without fine-tuning on UDC-VIT, or solely trained on UDC-VIT. Models without subscripts refer to those solely trained on UDC-VIT, meaning their PSNR, SSIM, and LPIPS values match those in Table 3 in the main body. The number of iterations represents the percentage of fine-tuning iterations relative to the total iterations in the original configurations provided by the authors.
$\text{DISCNet}_{s3}$ 16.81 0.7139 0.3293 UDC-SIT - UDC-VIT
DISCNet 24.70 0.8403 0.2675 UDC-VIT - UDC-VIT
[…]
UDC-UNet 28.00 0.8911 0.1779 UDC-VIT - UDC-VIT
Figure D.2: [Exp. 1] Comparison of restoration performance across different models when tested on the UDC-VIT dataset. (a) UDC-degraded (at the first row) and GT (at the second row) frames of the UDC-VIT dataset. Test frames from UDC-VIT by DISCNet (at the first row) and UDC-UNet (at the second row) trained on UDC-SIT (b) without fine-tuning on UDC-VIT, (c) with fine-tuning on UDC-VIT (10%), and (d) with fine-tuning on UDC-VIT (20%). (e) Test frames from UDC-VIT by DISCNet (at the first row) and UDC-UNet (at the second row), solely trained on UDC-VIT.
Experiment 2: impact of fine-tuning on VidUDC33K.
Table D.3: [Exp. 2] Restoration performance of DISCNet (Feng et al., 2021), UDC-UNet (Liu et al., 2022b), and DDRNet (Liu et al., 2024) when tested on VidUDC33K (Liu et al., 2024). They are trained on UDC-VIT, either with or without fine-tuning on VidUDC33K, or solely trained on VidUDC33K. Models without subscripts refer to those solely trained on VidUDC33K, meaning their PSNR, SSIM, and LPIPS values match those in Table 3 in the main body. The number of iterations represents the percentage of fine-tuning iterations relative to the total iterations in the original configurations provided by the authors.
$\text{DISCNet}_{s5}$ 18.73 0.7503 0.4159 UDC-VIT - VidUDC33K
$\text{DISCNet}_{s5z20}$ 28.89 0.9129 0.1727 UDC-VIT VidUDC33K ($10\%$) VidUDC33K
DISCNet 28.89 0.8405 0.2432 VidUDC33K - VidUDC33K
[…]
UDC-UNet 28.37 0.8361 0.2561 VidUDC33K - VidUDC33K
$\text{DDRNet}_{s5}$ 20.10 0.8313 0.3446 UDC-VIT - VidUDC33K
$\text{DDRNet}_{s5z20}$ 29.12 0.8994 0.2073 UDC-VIT VidUDC33K ($10\%$) VidUDC33K
DDRNet 31.91 0.9313 0.1306 VidUDC33K - VidUDC33K
[…]
Experiment 3: comparison of UDC-VIT and VidUDC33K.
This experiment evaluates the effect of fine-tuning on UDC-VIT when tested on UDC-VIT. We compare the performance of models trained on VidUDC33K, with or without fine-tuning on UDC-VIT, and models trained solely on UDC-VIT. For the models $\mathcal{M}$, we use DDRNet, the only publicly available pre-trained model trained on VidUDC33K. As shown in Table D.4, $\text{DDRNet}_{z20}$ without fine-tuning on UDC-VIT fails to adequately handle the complex, severe, and real-world degradations present in UDC-VIT. In contrast, $\text{DDRNet}_{z20s5}$, fine-tuned on UDC-VIT, demonstrates significant performance improvements over $\text{DDRNet}_{z20}$.
Table D.4: [Exp. 3] Restoration performance of DDRNet (Liu et al., 2024) trained on VidUDC33K (Liu et al., 2024), with and without additional fine-tuning on UDC-VIT. Models without subscripts refer to those trained solely on UDC-VIT, meaning their PSNR, SSIM, and LPIPS values match those in Table 3 in the main body. The number of iterations represents the percentage of fine-tuning iterations relative to the total iterations in the original configurations provided by the authors.
$\text{DDRNet}_{z20}$ 11.20 0.5331 0.5609 VidUDC33K - UDC-VIT
$\text{DDRNet}_{z20s5}$ 21.92 0.8302 0.2519 VidUDC33K UDC-VIT ($10\%$) UDC-VIT
DDRNet 24.68 0.8539 0.2218 UDC-VIT - UDC-VIT
Figure D.4: [Exp. 3] Restoration performance of DDRNet (Liu et al., 2024) tested on the UDC-VIT dataset. (a) UDC-degraded and (b) GT images from the UDC-VIT dataset. Restored images by (c) $\text{DDRNet}_{z20}$, (d) $\text{DDRNet}_{z20s5}$, and (e) DDRNet. The model $\text{DDRNet}_{z20}$ is pre-trained on VidUDC33K without fine-tuning on UDC-VIT, while $\text{DDRNet}_{z20s5}$ is pre-trained on VidUDC33K with fine-tuning on UDC-VIT, showing improved restoration performance. However, compared to the results in Figure D.3, the fine-tuned model still struggles to handle the real-world degradations present in the UDC-VIT dataset, since it is pre-trained on the synthetic VidUDC33K dataset.
However, as illustrated in Figure D.4, even with fine-tuning, $\text{DDRNet}_{z20s5}$ still shows limitations in handling specific real-world degradations, such as severe flares. It is important to note that models pre-trained on the real-world UDC-VIT dataset with fine-tuning perform well in restoring the synthetic VidUDC33K degradation in Experiment 2. On the other hand, models pre-trained on the synthetic VidUDC33K dataset with fine-tuning face challenges in effectively handling the real-world UDC-VIT degradation in Experiment 3. These findings highlight the importance of pre-training models on real-world datasets like UDC-VIT to capture better complex degradations that synthetic datasets cannot fully replicate.

### v2

@@ D.1 Dataset Acquisition
The alignment algorithm we use involves shifting, rotating, and cropping paired frames with DFT. The detailed algorithm is illustrated in Algorithm D.1. In this algorithm, following Ahn et al. [1]’s setting, we use $\lambda_{1}=\lambda_{3}=1$ and $\lambda_{2}=0$, and we do not apply rotation. Their experiments show that applying rotation reduces the Percentage of Correct Keypoints (PCK) when varying $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\theta_{\text{rotation}}$.
Figure D.1: Frequency analysis based on the conceptual illustration for paired frames involving shifts without degradation. (a) The original frame $\mathcal{G}$ consists of multiple sinusoidal gratings. The inverse $DFT$ applied to $\mathcal{F}_{G}(u,v)$ produces each sinusoidal grating. (b) The phase difference between $\mathcal{G}$ and $\mathcal{D}$. (c) The spatially shifted frame $\mathcal{D}$ in the spatial domain comprises multiple sinusoidal gratings, as in (a). (d) The amplitude difference between $\mathcal{G}$ and $\mathcal{D}$, showing no difference.
The loss function in Eq. 1 in the main body enables the incorporation of both local (i.e., MSE) and global (i.e., DFT) information across spatial and frequency domains. Using DFT to align the paired frames offers a significant advantage because it can decompose a frame into its constituent spatial frequency components. Fig. D.1(a) and (c) depict paired frames $\mathcal{G}$ and $\mathcal{D}$ comprising multiple sinusoidal gratings, indicating a noticeable spatial shift. Fig. D.1(b) and (d) represent the phase and amplitude differences between the paired frames, respectively. Thus, reducing the phase component is critical for effectively aligning the paired frames for the same scene.
Algorithm D.1 Alignment of paired images $I_{G}$ and $I_{D}$ [1].
[…]
@@
[…]
@@
[…]
@@
D.2 Dataset Details and Statistics
@@
From a pool of 647 videos, we have randomly selected 510 for training, 69 for validation, and 68 for the test set. The UDC-VIT dataset is available in PNG format accompanied by a conversion script from PNG to NPY. We have annotated each video pair, providing a detailed overview of the total count and the distribution of different annotation labels. The video pairs are thoughtfully categorized into various conditions, including the presence of flare and light sources, the presence of humans and their motion types, and whether the scene is indoor or outdoor.
Figure D.2: The dataset distribution. The parenthesis beside a label is the encoding of the label. Note that a video pair can have multiple annotation labels. The distribution of (a) lighting conditions, (b) human presence and their actions, and (c) shooting location.
Table D.1: Prescribed instructions from IRB-approved participant information sheet. Out of thirty shots per person, videos displaying issues such as being out-of-focus are eliminated from the dataset.
@@
A. Participants will undergo a 30-shot photo session using UDC and regular cameras with the following motions:
@@
Since the UDC is located beneath the display and operates in low-light environments, it is necessary to capture data in various locations (indoors and outdoors) and lighting conditions (bright and dark) to account for the diverse degradation patterns of the UDC. Additionally, recognizing individuals from different angles is crucial for tasks such as face recognition, especially in financial authentication. This requires deep-learning models capable of restoring a subject’s appearance from multiple perspectives (e.g., front, left, and right), and thus, a dataset with multi-angle captures is essential. The recorded videos will be publicly released as a dataset to support UDC research.
@@
A. As a token of appreciation, participants who complete all 30 video captures will receive a \$40 Starbucks gift card. However, participants who withdraw before completing all 30 shots or request video deletion will not be compensated. If a participant requests video deletion after receiving compensation, they must return the full amount.
@@
We have obtained the approval of the Institutional Review Board (IRB) for the UDC-VIT dataset, as our research involves human subjects. This rigorous process ensures the highest standards of research ethics. Using IRB-approved procedures, we enlisted 22 voluntary research participants. As shown in Tab. D.1, the IRB-approved information sheet provides detailed instructions to the participants.
Similarly, it is essential to note that the users of the UDC-VIT dataset are engaged in research involving human subjects. Therefore, users are required to obtain IRB approval by the regulations of their respective countries. When users download the dataset, instructions on IRB approval will be provided, as shown in Fig. D.3.
D.3 Rigorous Maintenance Plan
This section provides easy accessibility and a rigorous maintenance plan for the UDC-VIT dataset.
@@ D.3 Rigorous Maintenance Plan
The UDC-VIT dataset will be publicly available on our research group’s website, as shown in Fig. D.3. Users can request access by completing a form on the site; a download link will be sent via email upon submission. Detailed instructions are available at https://github.com/mcrl/UDC-VIT. Hosting the dataset on the research group’s website ensures long-term availability, while handling contact and bug reports via email facilitates ongoing maintenance and updates.
Figure D.3: Our research group’s official website providing detailed information and a download link for the UDC-VIT dataset.
@@
The UDC-VIT dataset is licensed under Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International (CC BY-NC-SA 4.0). Under this license, users of the UDC-VIT dataset can freely utilize, share, and modify this work by adequately attributing the original author, distributing any derived works under the same license, and utilizing it exclusively for non-commercial purposes. It is essential to mention that the UDC-VIT dataset is restricted to UDC research purposes only, as outlined in our IRB documentation. Detailed information on this license can be found on the official Creative Commons website.